OpenStudy (logan13):

Which ordered pair is the solution to the system of equations? A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3)

OpenStudy (welshfella):

to solve this multiply the second equation by 4 and add this new equation to the first one. this will eliminate y so you can find x Then find x by plugging in the value oy y into equation 1.

OpenStudy (logan13):

so 1 -9

OpenStudy (welshfella):

*Then find y by plugging in the value of x into equation 1.

OpenStudy (welshfella):

No

OpenStudy (logan13):

@welshfella a

OpenStudy (welshfella):

lets do the thing x - 4y = 6 3x + y = -8 * this by 4 12x + 4y = -32 x - 4y = 6 ADD:- 13x = -26 now divide both sides by 13 to get the value of x I think you can finish this off.

OpenStudy (logan13):

@welshfella so D

OpenStudy (welshfella):

you've only got to work out x from 13x = -26 and you'll see which option it is straight away.